# Sloth: scaling laws for LLM skills to predict multi-benchmark performance across families

### v4

@@ 4 Sloth in practice
In this section, we present experimental results that provide evidence of the usefulness of Sloth. We perform experiments on a set of twelve benchmarks and state-of-the-art LLM families, including LLaMa 3 (Dubey et al., 2024), Qwen 2 (Yang et al., 2024), and Yi 1.5 (Young et al., 2024). We explore the following applications: (i) benchmark performance prediction for larger models from a specific LLM family, (ii) interpretability of the scaling of skills (can help practitioners allocate resources based on the skills of interest), and (iii) downstream tasks performance prediction.
@@ 4.1 Data
We expand the dataset made available by Ruan et al. (2024), including more models from the HuggingFace Open LLM leaderboard v1 (Beeching et al., 2023) and v2 (Fourrier et al., 2024). In our extended dataset, we have a total of 30 families [^1] If we consider that instruct and base models are from different families, we end up with 53 families., which 28 are on v1 of the Open LLM Leaderboard and 17 families measured on v2 of the Open LLM Leaderboard. Furthermore, there are 15 families at the intersection of the two versions. Furthermore, we collect data and present results on the performance of a variety of instruction-tuned versions of the base models we consider. As far as we are aware, our dataset is the most comprehensive among prior works on benchmark data scaling laws. Please check Appendix G for details on the included models.
@@ 4.2 Comparing scaling laws in terms of prediction errors
In this section, we compare the predictive power of different scaling laws in predicting LLM performance in all the considered benchmarks; we focus on the two versions of the Open LLM Leaderboard, which include 12 benchmarks: GSM8k (Cobbe et al., 2021), MATH Lvl 5 (Hendrycks et al., 2021), MMLU (Hendrycks et al., 2020), MMLU-PRO (Wang et al., 2024), BBH (Suzgun et al., 2022), GPQA (Rein et al., 2023), MUSR (Sprague et al., 2023), TruthfulQA (Lin et al., 2021), HellaSwag (Zellers et al., 2019), Winogrande (Sakaguchi et al., 2019), ARC (Clark et al., 2018), and IFEval (Zhou et al., 2023). We apply a leave-one-out cross-validation algorithm to obtain test errors for each family of models. We consider base models and instruct models to belong to distinct families (they will not share the same intercept in our model, for example), but we do not include the instruct (resp. base) family in the training set when the corresponding base (resp. instruct) family is in the test set. Moreover, we do not test older versions of recent families if they are available in the training set, e.g., we do not include LLaMa 2 in the set of test families if LLaMa 3 is present in the training set. In this section, we present results for the two leaderboards separately; in Figures 11 and 16 of the Appendix, we also present results for the intersection of the two leaderboards.
@@
In the first set of experiments, we consider the case in which only the smallest model of the test family is observed at training time. Because of that reason, we cannot fit the general scaling law in (2.2) in which both the intercept and slope are family dependent. In this scenario, we consider our main baselines to be (i) the model in (2.2) with shared intercept and slope (Owen, 2024) (“FLOPs (shared intercept)”), (ii) the same model but only with shared slope (“FLOPs”), (iii) a version of the PCA idea [^2] We include more details about this approach in Appendix F. proposed by Ruan et al. (2024) in which we predict the principal components using the FLOPs model with shared slope that are then mapped to the benchmark scores (“PCA + FLOPs”), (iv) and our model with trainable activation function but assuming $\Lambda$ is identity (“Size and Tokens”; implies $d=J$). Moreover, we include two versions of Sloth. In the “basic” one, we fix $\sigma$ to be sigmoid, and $\gamma_{j}$’s are given by the $100\%$ over the number of alternatives in the case of multiple-choice benchmarks [^3] When the benchmark has subsections with a different number of alternatives, we compute the asymptote parameters per subsection and then compute an overall asymptote using a weighted average in which the weights are proportional to the number of examples in each subsection. and 0 otherwise, except for TruthfulQA, which we compute the first percentile of the scores coming from the full Open LLM Leaderboard and fix the lower asymptote to that value.
@@
Figure 1 gives the results for the first set of experiments. It depicts the average mean absolute error of all methods when predicting LLM benchmark performance, which is measured in percentage points. It shows the competitiveness of Sloth in terms of prediction quality. One important thing to notice is that Sloth errors are similar or lower than the “Size and Tokens” variant, suggesting that the assumed low-dimensional structure between benchmarks results is a good inductive bias. We highlight that the analysis includes recent families like LLaMa 3, Qwen 2, and Yi 1.5. For more details on the tested models and extra related results, including model-specific results, please check Appendix H.1. The extra results are qualitatively similar to the ones in Figure 1, in which Sloth often beats the baselines. In a second set of experiments, we consider the case in which the two smallest models of the test family are observed at training time. In this way, we can fit (2.2) making both parameters family dependent. The results are qualitatively similar to the one presented on Figure 1 and are presented in Appendix H.2.
@@
In an extra set of experiments, we show that family-specific intercept models are not always needed; we can still get good prediction results for some benchmarks even if we consider a shared intercept between families. The advantage of this approach is that we can claim for a general scaling law that holds for all families. Figure 3 shows us a subset [^4] We selected the best $d$ for both versions of Sloth. of Figure 11 in the appendix and it is built under the same conditions as Figure 1. It is possible to see that, for a subset of benchmarks Sloth with shared intercept is a strong alternative to the FLOPs model used by Owen (2024). In some cases, it gets similar prediction errors relative to more complete versions of Sloth.
@@
Figure 5: We compare the skills of base (x-axis) and instruction-tuned models (y-axis); if a model lies in the 45-degree line, it means that the model has the same skill level in its base and instruct version. Gains from instruction tuning (IT) for different families on three latent skills. Findings include a large and positive impact on “Instruction Following” and that provide much larger variations in this skill when compared to inputs seen in Figure 4. Moreover, IT had a moderate and negative effect on “Reasoning” and mixed effects on “Knowledge”.
@@ 4.3 Interpreting the latent skills
In this section, we use the intersection between the two leaderboards, aiming to get more insights from the combined data. Since we have an identifiability result for the “basic” version of Sloth, in which we fix the lower asymptotes $\gamma_{j}$’s and the link function to be sigmoid (see Section B), we opt for interpreting that version of the model. We set $d=3$ as that model version achieved the best prediction results in Figure 11. Figure 2 illustrates the model loadings, $\Lambda$, from which we assign names to the three dimensions based on our subjective interpretation. We include the loadings for $d=2$ and $d=4$ in Appendix I. To complement our exploration, we include Figure 4, which gives us the level curves of different skills (disregarding the family-specific intercept term), and Figure 5 that compare the skills of base and instruction-tuned models; in this figure, we include LLM families with more number of models. In both figures, the numbers are given in terms of standard deviations as the skills are standardized to have zero mean and unitary standard deviation.
@@
Reasoning skill: The first dimension, with strong loadings from benchmarks such as GSM8K, MATH, GPQA, MMLU(-PRO), BBH, and MUSR, is labeled “Reasoning.” The benchmarks GSM8k and MATH Lvl 5 consist entirely of mathematical word problems while MMLU/MMLU-PRO and GPQA also contain mathematical and advanced science questions. On the other hand, BBH includes logic deduction and linguistic reasoning puzzles. The strong dependence of BBH on the “Reasoning” skill suggests that in language models, there is an association between logical reasoning, general linguistic ability, and mathematical ability. Finally, MuSR is a benchmark that evaluates “multistep soft reasoning tasks specified in a natural language narrative” (Sprague et al., 2023). Figure 4 shows that Reasoning is primarily a function of model size, with a small dependence on the number of training tokens used. Moreover, the first plot of Figure 5 compares base models versus their instruction-tuned versions in terms of Reasoning and we found that there is no clear rule: instruction tuning can either increase or decrease the ability of an LLM to reason. These findings are robust for different values of $d$ as we can see in the figures of Appendix I.
@@
Knowledge skill: The second dimension is positively loaded on ARC, HellaSwag, and Winogrande. These three benchmarks measure the ability of LLMs to remember common sense and basic knowledge; we denominate this skill as “Knowledge”. More specifically, ARC consists of grade school-level science questions, HellaSwag is meant for sentence completion for common scenarios, and Winogrande common sense pronoun resolution problems. Contrasting with Reasoning, Figure 4 shows that Knowledge is highly influenced by both model size and number of training tokens. Moreover, we can see that the range of standard deviations in the middle plot is much greater than in the other two plots, giving us evidence that this skill might be more sensitive to increases in compute resources and less dependent on the LLM families themselves. On the other hand, Figure 5 does not show any strong evidence of the effect of instruction tuning on this skill. These findings are similar to the ones reported in Appendix I for different values of $d$.
@@
Instruction following skill: IFEval, which is positively and heavily loaded in this skill, measures how well language models produce answers that follow a verifiable set of instructions; for example, including a keyword $x$ number of times in responses. Therefore, we call it “Instruction Following”. An interesting fact is that instruction tuning has a strong positive effect on this skill for all depicted families we can see in Figure 5. The effect can also be observed in Figure 27 of the appendix. When $d=2$, instruction following gets mixed with other skills and we are not able to see this effect. Regarding Figure 4, we see that Instruction Following depends on both model size and tokens. Unfortunately, this interpretation does not hold when $d=4$ as seen in Appendix I.
@@
4.4 Predicting LLM performance on downstream tasks
Another useful application of Sloth, which is inspired by Ruan et al. (2024), is to predict the performance in a downstream task for a large model from a relatively small number of prior performance observations from that task. We use Sloth to estimate the latent skills of hypothetical LLMs and then use them to predict the performance of those LLMs in downstream tasks. With this approach, we expand on the experiments of Ruan et al. (2024), which do not consider performance prediction of hypothetical LLMs.
@@ 4.4 Predicting LLM performance on downstream tasks
We evaluate this pipeline on two downstream tasks, predicting the performance of meta-llama-3-70B and meta-llama-3-70B-instruct on code completion and meta-llama-3-70B-instruct on emotional intelligence tasks. We fit the same model shown in Section 4.3, but do not include meta-llama-3-70B or meta-llama-3-70B-instruct in the training set (see Figure 31 for the loadings of the latent skills, which is similar to Figure 2). Next, using either HumanEval (Chen et al., 2021) or EQ bench data (Paech, 2024), we fit a regression model with logistic link using latent skills as features and performance on the downstream task as target. Together, this provides us with sufficient information to predict the performance of the held-out models on both tasks with decent accuracy. Figure 6 depicts this logistic curve and the actual values for HumanEval; EQ bench results can be found in Appendix J. Moreover, we can see that “Reasoning” is by far the most important skill in predicting coding ability while a mixture of “Reasoning” and “Knowledge” is needed for emotional intelligence (see Figure 31 for a more accurate interpretation of the loadings).
In Appendix J, a similar test is provided for agentic capability measured by AgentBench (Liu et al., 2023), although to avoid overfitting, in this case, we must fit Sloth with no family-specific intercept.
@@ 4.5 Predicting performance behavior with scaled inference compute
In this section, we demonstrate how our method, Sloth, combined with concepts like Item Response Theory (IRT) (Reckase, 2009), can predict how the performance of an LLM scales with increased inference through repeated sampling (Brown et al., 2024). For this experiment, we utilize the MATH dataset (Hendrycks et al., 2021), made available by Brown et al. (2024), which evaluates 10 LLMs also included in our dataset. These models are part of the LLaMa 3 Instruct, Gemma, and Pythia families. The process is:
@@
    Fit a logistic regression model for each MATH question in Brown et al. (2024)’s data, using the skills $\theta_{i}(s,t)$ of the 7 training LLMs as covariates to predict the probability of correctly solving those problems.
@@
Figure 7 illustrates that Sloth can accurately predict test-time scaling behavior for these models. Note that unlike scaling law in Brown et al. (2024), Sloth can predict inference compute gains for hypothetical LLMs before committing resources to training them. This highlights a practical application where practitioners can estimate the potential performance of a hypothetical model at test time given a computational budget.
Figure 7: Sloth can accurately predict test-time scaling behavior for LLaMa 3 Instruct, Gemma, and Pythia models when coupled with Item Response Theory (IRT).

### v5

@@ 4 Sloth in practice
In this section, we present experimental results that provide evidence of the usefulness of Sloth. We perform experiments on a set of twelve benchmarks and state-of-the-art LLM families, including LLaMa 3 (dubey2024llama), Qwen 2 (yang2024qwen2), and Yi 1.5 (young2024yi). We explore the following applications: (i) benchmark performance prediction for larger models from a specific LLM family, (ii) interpretability of the scaling of skills (can help practitioners allocate resources based on the skills of interest), and (iii) downstream tasks performance prediction.
@@ 4.1 Data
We expand the dataset made available by ruan2024observational, including more models from the HuggingFace Open LLM leaderboard v1 (openllmldrboard) and v2 (open-llm-leaderboard-v2). In our extended dataset, we have a total of 30 families [^2] If we consider that instruct and base models are from different families, we end up with 53 families., which 28 are on v1 of the Open LLM Leaderboard and 17 families measured on v2 of the Open LLM Leaderboard. Furthermore, there are 15 families at the intersection of the two versions. Furthermore, we collect data and present results on the performance of a variety of instruction-tuned versions of the base models we consider. As far as we are aware, our dataset is the most comprehensive among prior works on benchmark data scaling laws. Please check Appendix G for details on the included models.
@@ 4.2 Comparing scaling laws in terms of prediction errors
In this section, we compare the predictive power of different scaling laws in predicting LLM performance in all the considered benchmarks; we focus on the two versions of the Open LLM Leaderboard, which include 12 benchmarks: GSM8k (cobbe2021trainingverifierssolvemath), MATH Lvl 5 (hendrycks2021measuringmathematicalproblemsolving), MMLU (hendrycks2020Measuring), MMLU-PRO (wang2024mmluprorobustchallengingmultitask), BBH (suzgun2022challengingbigbenchtaskschainofthought), GPQA (rein2023gpqagraduatelevelgoogleproofqa), MUSR (sprague2023musr), TruthfulQA (lin2021truthfulqa), HellaSwag (zellers2019hellaswagmachinereallyfinish), Winogrande (sakaguchi2019winograndeadversarialwinogradschema), ARC (clark2018thinksolvedquestionanswering), and IFEval (zhou2023instructionfollowingevaluationlargelanguage). We apply a leave-one-out cross-validation algorithm to obtain test errors for each family of models. We consider base models and instruct models to belong to distinct families (they will not share the same intercept in our model, for example), but we do not include the instruct (resp. base) family in the training set when the corresponding base (resp. instruct) family is in the test set. Moreover, we do not test older versions of recent families if they are available in the training set, e.g., we do not include LLaMa 2 in the set of test families if LLaMa 3 is present in the training set. In this section, we present results for the two leaderboards separately; in Figures 12 and 17 of the Appendix, we also present results for the intersection of the two leaderboards.
@@
In the first set of experiments, we consider the case in which only the smallest model of the test family is observed at training time. Because of that reason, we cannot fit the general scaling law in (2.2) in which both the intercept and slope are family dependent. In this scenario, we consider our main baselines to be (i) the model in (2.2) with shared intercept and slope (owen2024predictable) (“FLOPs (shared intercept)”), (ii) the same model but only with shared slope (“FLOPs”), (iii) a version of the PCA idea [^3] We include more details about this approach in Appendix E. proposed by ruan2024observational in which we predict the principal components using the FLOPs model with shared slope that are then mapped to the benchmark scores (“PCA + FLOPs”), (iv) and our model with trainable activation function but assuming $\Lambda$ is identity (“Size and Tokens”; implies $d=J$). Moreover, we include two versions of Sloth. In the “basic” one, we fix $\sigma$ to be sigmoid, and $\gamma_{j}$’s are given by the $100\%$ over the number of alternatives in the case of multiple-choice benchmarks [^4] When the benchmark has subsections with a different number of alternatives, we compute the asymptote parameters per subsection and then compute an overall asymptote using a weighted average in which the weights are proportional to the number of examples in each subsection. and 0 otherwise, except for TruthfulQA, which we compute the first percentile of the scores coming from the full Open LLM Leaderboard and fix the lower asymptote to that value. In previous sections, we mentioned that Sloth is parameter efficient; we include a parameter count analysis in Appendix F where we compare our model with other well-performing baselines.
@@
Figure 1 gives the results for the first set of experiments. It depicts the average mean absolute error of all methods when predicting LLM benchmark performance, which is measured in percentage points. It shows the competitiveness of Sloth in terms of prediction quality. One important thing to notice is that Sloth errors are similar or lower than the “Size and Tokens” variant, suggesting that the assumed low-dimensional structure between benchmarks results is a good inductive bias. We highlight that the analysis includes recent families like LLaMa 3, Qwen 2, and Yi 1.5. For more details on the tested models and extra related results, including model-specific results, please check Appendix H.2. The extra results are qualitatively similar to the ones in Figure 1, in which Sloth often beats the baselines. In a second set of experiments, we consider the case in which the two smallest models of the test family are observed at training time. In this way, we can fit (2.2) making both parameters family dependent. The results are qualitatively similar to the one presented on Figure 1 and are presented in Appendix H.3. Moreover, we include Figure 10 in the appendix, which is a version of Figure 1 in which Mean Absolute Percentage Error (MAPE) is used instead of MAE; results are qualitatively similar.
In an extra set of experiments, we show that family-specific intercept models are not always needed; we can still get good prediction results for some benchmarks even if we consider a shared intercept between families. The advantage of this approach is that we can claim for a general scaling law that holds for all families. Figure 3 shows us a subset [^5] We selected the best $d$ for both versions of Sloth. of Figure 12 in the appendix and it is built under the same conditions as Figure 1. It is possible to see that, for a subset of benchmarks Sloth with shared intercept is a strong alternative to the FLOPs model used by owen2024predictable. In some cases, it gets similar prediction errors relative to more complete versions of Sloth.
@@
Figure 5: We compare the skills of base (x-axis) and instruction-tuned models (y-axis); if a model lies on the 45-degree line, it means that the model has the same skill level in its base and instruct versions. Gains from instruction tuning (IT) for different families on three latent skills. Findings include a large and positive impact on “Instruction Following” and that provides much larger variations in this skill when compared to inputs seen in Figure 4. Moreover, IT had a moderate and negative effect on “Reasoning” and mixed effects on “Knowledge”.
@@ 4.3 Interpreting the latent skills
In this section, we use the intersection between the two leaderboards, aiming to get more insights from the combined data. Since we have an identifiability result for the “basic” version of Sloth, in which we fix the lower asymptotes $\gamma_{j}$’s and the link function to be sigmoid (see Section A), we opt for interpreting that version of the model. We set $d=3$ as that model version achieved the best prediction results in Figure 12. Figure 2 illustrates the model loadings, $\Lambda$, from which we assign names to the three dimensions based on our subjective interpretation. We include the loadings for $d=2$ and $d=4$ in Appendix I. To complement our exploration, we include Figure 4, which gives us the level curves of different skills (disregarding the family-specific intercept term), and Figure 5 in the appendix that compares the skills of base and instruction-tuned models; in this figure, we include LLM families with more number of models. In both figures, the numbers are given in terms of standard deviations as the skills are standardized to have a zero mean and unitary standard deviation.
@@
Reasoning skill: The first dimension, with strong loadings from benchmarks such as GSM8K, MATH, GPQA, MMLU(-PRO), BBH, and MUSR, is labeled “Reasoning.” The benchmarks GSM8k and MATH Lvl 5 consist entirely of mathematical word problems while MMLU/MMLU-PRO and GPQA also contain mathematical and advanced science questions. On the other hand, BBH includes logic deduction and linguistic reasoning puzzles. The strong dependence of BBH on the “Reasoning” skill suggests that in language models, there is an association between logical reasoning, general linguistic ability, and mathematical ability. Finally, MuSR is a benchmark that evaluates “multistep soft reasoning tasks specified in a natural language narrative” (sprague2023musr). Figure 4 shows that Reasoning is primarily a function of model size, with a small dependence on the number of training tokens used. Moreover, the first plot of Figure 5 in the appendix compares base models versus their instruction-tuned versions in terms of Reasoning, and we found that there is no clear rule: instruction tuning can either increase or decrease the ability of an LLM to reason. These findings are robust for different values of $d$ as we can see in the figures of Appendix I.
@@
Knowledge skill: The second dimension is positively loaded on ARC, HellaSwag, and Winogrande. These three benchmarks measure the ability of LLMs to remember common sense and basic knowledge; we denominate this skill as “Knowledge”. More specifically, ARC consists of grade school-level science questions, HellaSwag is meant for sentence completion for common scenarios, and Winogrande common sense pronoun resolution problems. Contrasting with Reasoning, Figure 4 shows that Knowledge is highly influenced by both model size and number of training tokens. Moreover, we can see that the range of standard deviations in the middle plot is much greater than in the other two plots, giving us evidence that this skill might be more sensitive to increases in compute resources and less dependent on the LLM families themselves. On the other hand, Figure 5 in the appendix does not show any strong evidence of the effect of instruction tuning on this skill. These findings are similar to the ones reported in Appendix I for different values of $d$.
@@
Instruction following skill: IFEval, which is positively and heavily loaded in this skill, measures how well language models produce answers that follow a verifiable set of instructions; for example, including a keyword $x$ number of times in responses. Therefore, we call it “Instruction Following”. An interesting fact is that instruction tuning has a strong positive effect on this skill for all depicted families we can see in Figure 5 in the appendix. The effect can also be observed in Figure 28 of the appendix. When $d=2$, instruction following gets mixed with other skills and we are not able to see this effect. Regarding Figure 4, we see that Instruction Following depends on both model size and tokens. Unfortunately, this interpretation does not hold when $d=4$ as seen in Appendix I.
4.4 Predicting LLM performance on downstream tasks
Another useful application of Sloth, which is inspired by ruan2024observational, is to predict the performance in a downstream task for a large model from a relatively small number of prior performance observations from that task. We use Sloth to estimate the latent skills of hypothetical LLMs and then use them to predict the performance of those LLMs in downstream tasks. With this approach, we expand on the experiments of ruan2024observational, which do not consider performance prediction of hypothetical LLMs.
@@ 4.4 Predicting LLM performance on downstream tasks
We evaluate this pipeline on two downstream tasks, predicting the performance of meta-llama-3-70B and meta-llama-3-70B-instruct on code completion and meta-llama-3-70B-instruct on emotional intelligence tasks. We fit the same model shown in Section 4.3, but do not include meta-llama-3-70B or meta-llama-3-70B-instruct in the training set (see Figure 32 for the loadings of the latent skills, which is similar to Figure 2). Next, using either HumanEval (chen2021codex) or EQ bench data (paech2024eqbenchemotionalintelligencebenchmark), we fit a regression model with logistic link using latent skills as features and performance on the downstream task as target. Together, this provides us with sufficient information to predict the performance of the held-out models on both tasks with decent accuracy. Figure 6 depicts this logistic curve and the actual values for HumanEval; EQ bench results can be found in Appendix J. Moreover, we can see that “Reasoning” is by far the most important skill in predicting coding ability while a mixture of “Reasoning” and “Knowledge” is needed for emotional intelligence (see Figure 32 for a more accurate interpretation of the loadings).
In Appendix J, a similar test is provided for agentic capability measured by AgentBench (liu2023AgentBench), although to avoid overfitting, in this case, we must fit Sloth with no family-specific intercept.
@@ 4.5 Predicting performance behavior with scaled inference compute
In this section, we demonstrate how our method, Sloth, combined with concepts like Item Response Theory (IRT) (reckase2009multidimensional), can predict how the performance of an LLM scales with increased inference through repeated sampling (brown2024large). For this experiment, we utilize the MATH dataset (hendrycks2021measuringmathematicalproblemsolving), made available by brown2024large, which evaluates 10 LLMs also included in our dataset. These models are part of the LLaMa 3 Instruct, Gemma, and Pythia families. The process is:
@@
    Fit a logistic regression model for each MATH question in brown2024large’s data, using the skills $\theta_{i}(s,t)$ of the 7 training LLMs as covariates to predict the probability of correctly solving those problems.
@@
Figure 7: Sloth can accurately predict test-time scaling behavior for LLaMa 3 Instruct, Gemma, and Pythia models for MATH performance when coupled with Item Response Theory.
Figure 7 illustrates that Sloth can accurately predict test-time scaling behavior for these models. Note that unlike scaling law in brown2024large, Sloth can predict inference compute gains for hypothetical LLMs before committing resources to training them. This highlights a practical application where practitioners can estimate the potential performance of a hypothetical model at test time, given a computational budget.
4.6 Compute-optimal scaling of skills
One relevant question is: given a certain budget in FLOPs, how do we invest it to maximize one skill of interest? This type of analysis is novel for skills and it has only been carried out for validation loss (e.g., by kaplan2020scalinglawsneurallanguage; hoffmann2022training). A summary of the mathematical formulation of this question is given in the following and exposed in detail in Appendix B. For each model family $i$ and skill $k$, we have
$\theta_{ik}(s,t)=\alpha_{ik}+\beta_{k0}\log(s)+\beta_{k1}\log(t)+\beta_{k2}\log(s)\log(t).$
Define $u=\log(s)$, $v=\log(t)$, and $l=\log(c)-\log(6)$. For simplicity, here we consider a simple, but widely used (kaplan2020scalinglawsneurallanguage; hoffmann2022training), compute budget constraint $6st=c$, which is equivalent to $u+v=l$. We add extra constraints on $u$ and $v$ based on the training data support to avoid unreasonable out-of-support predictions. We optimize
$\max_{u,v}\alpha_{ik}+\beta_{k0}u+\beta_{k1}v+\beta_{k2}uv\quad\text{subject to}\quad u+v=l,\quad u\in[\underline{u},\overline{u}],\;v\in[\underline{v},\overline{v}].$
FLOPs (1e19) Instruction Following
Params (B) Tokens (T)
100 0.16 1.04
578 0.30 3.24
3346 0.72 7.78
19360 2.15 15.0
112005 12.44 15.0
648000 72.0 15.0
Table 1: Optimal parameter and token allocation.
Substituting v=luv=l-u reduces this to a quadratic function $g_{ik}(u)$, whose maximizer, within the observed range $\mathcal{U}=\left[\max(l-\overline{v},\underline{u}),\min(l-\underline{v},\overline{u})\right]$, determines the optimal allocation. Our analysis indicates that the optimal values for $s$ and $t$ do not depend on the model family. Table 1 reports our results for instruction following; a table with results for the other two skills is shown in Appendix B.